OpenStudy (anonymous):

needing help to slove this: (5y^5-y^2) divided by -y

OpenStudy (hba):

I am sorry i was busy.

OpenStudy (anonymous):

it's cool lol

OpenStudy (hba):

Lets start solving this then.

OpenStudy (anonymous):

ok lol its really confusing to me and i need to pass my test or i will fail the whole thing and have to redo it all over again :((

OpenStudy (hba):

\[\frac{ 5y^5-5y^2 }{ -y }\] This is your question right ?

OpenStudy (anonymous):

yupp :)) lol

OpenStudy (hba):

\[\frac{ 5y^5 }{ -y }+\frac{- 5y^2 }{ -y }\] Do you understand this ?

OpenStudy (anonymous):

Not really :/

OpenStudy (hba):

I just seprated them.

OpenStudy (anonymous):

ok lol sorry was at lunch